# Practicalities of State-Dependent and Threshold Delay Differential Equations

### v1

@@ 10 IVPs with Distributed and Threshold Delays
While it is possible to develop solvers for specific problems (see Langlois et al. (2017); Cassidy et al. (2022)), it is a lot of work to reinvent and recode all the functionality of a sophisticated IVP solver for each new problem. Until a comprehensive numerical RFDE IVP solver is available, it is convenient and time saving to leverage the existing discrete delay software packages to solve to solve problems with distributed delays.
@@ 11 Numerical Continuation and Bifurcation Detection
To apply DDE-Biftool to (2.2),(2.1) we introduce extra dummy delays to discretize the the threshold integral. Recall (2.4), which implies that the delay $\tau=\tau(u_{t})$ satisfies
@@
We discretize the time interval $t-\tau_{\textit{max}},t]$ uniformly with a sequence of mesh points
$t=x_{0}>x_{1}>...>x_{N}=t-\tau_{\textit{max}},$
@@
It remains to find $\theta\in[0,1)$ so that $J_{h}(j+\theta)=a$ and and hence $\tau_{j}=(j+\theta)a/NV_{\textit{min}}$. To do this we extent the definition of $J_{h}(j)$ to non-integer values of $j$ as illustrated in Figure 4.6. See Gedeon et al. (2022); Wang and Humphries (2022) for further details.
@@ 12 Scalar Threshold Delay Example
Figure 5.1: Bifurcation diagram of (2.2)-(2.1) with $g$ and $V$ both increasing and parameters $\beta=1.4$, $\mu=0.2$, $g^{-}=0.5$, $g^{+}=1$, $\gamma=0.8$, $\theta_{v}=0.5$, $\theta_{g}=a=1$, $v^{-}=0.1$ and $v^{+}=2$. (a) One parameter continuation in $\gamma$ with $m=n=20$. (b) Two-parameter continuations in $m=n$ and $\gamma$ of the fold and the Hopf bifurcations. (Reproduced from Gedeon et al. (2025) under $C$-BY 4.0 license)
When $g(u)$ and $V(u)$ are both increasing functions, it is possible to have up to five coexisting steady states as illustrated in Figure 5.1(a). In particular, a region of tristability is observed, characterized by three stable steady states separated by two unstable ones. This contrasts sharply with the dynamics of the same operon model with constant state delay, for which at most three steady states are observed when $g$ is an increasing Hill function. If we flip the axes in the Figure 5.1(a) we see that each value of $u>0$ is a steady state for a unique value of $\gamma$, and that the values of $\theta_{g}$ and $\theta_{v}$ organise the locations of the fold bifurcations.
@@
Figure 5.2: One parameter bifurcation diagrams of (2.2)-(2.1) with $g$ constant, and $V$ an increasing function defined by (5.1) with (a) $m=50$ (b) $m=2$ (c) $m=2.25$ and (d) $m=2.35$. Remaining parameters: $\beta=1.4$, $\mu=0.2$, $g^{-}=g^{+}=1$, $\theta_{v}=1$, $a=1$, $v^{-}=0.1$ and $v^{+}=2$. In (b) periodic orbits are represented by the maximum and minimum values of $u$ along the periodic solution profile, and in the other panels by the $L^{2}$-norm (5.2). (Reproduced from Gedeon et al. (2025) under $C$-BY 4.0 license)
@@
In contrast, in Figure 5.2(b) plot both $\max u(t)$ and $\min u(t)$ to represent the periodic orbit, which clearly shows the amplitude of the solution. Both representations can be useful on bifurcation diagrams, since all three expressions are equal to the steady state value at a Hopf bifurcation, and thus periodic orbits are seen to emanate from a steady states at a Hopf bifurcation. However, the $L^{2}$-norm (5.2) has the extra property that $\|u\|_{2}\to u_{s}$ as the solution converges to a homoclinic orbit to a saddle steady state $u_{s}$, as seen in both Figures 5.1 and 5.2.
@@ 13 Two Linearly State-Dependent Delays
Figure 5.3: Bifurcation diagrams of (1.22) as $\kappa_{1}$ is varied showing the amplitude and stability of periodic orbits with (a) $\kappa_{2}=1$ (b) $\kappa_{2}=2.3$. In (b) the grey curves denote stable invariant tori. (© AIMS; Reproduced from Humphries et al. (2012) with permission)
@@
Figure 5.4: Poincaré section $u_{t}(0)=0$ of the quasi periodic torus from Figure 1.5 generated from a single solution of (1.22) with $\kappa_{1}=4.44$ and $\kappa_{2}=3$. (a) Solution segments $u_{t}$ on the Poincaré section. (b) Projection of the solution onto $(u(t),u(t-a_{1}),u(t-a_{2}))$-space (light blue), together with the two-dimensional projection of the Poincaré section. (©SIAM; Reproduced from Calleja et al. (2017) with permission)
Solutions oscillate about the steady state $u^{*}=0$, so a natural Poincaré section to study the dynamics is $\bigl\{\varphi\in C:\varphi(0)=0\;\&\;\varphi^{\prime}(0)<0\bigr\}$. For finite dimensional dynamical systems the Poincaré section reduces the dimension by one, but $C$ is infinite dimensional, resulting in an infinite dimensional Poincaré section. Plotting function segments on the Poincaré section with $u_{t}(0)=0$ is not very revealing, as seen in Figure 5.4(a).
@@
Figure 5.5: Two-parameter continuation of Hopf (blue), torus (red) and fold (brown) bifurcations in the $(\kappa_{1},\kappa_{2})$-parameter plane show many resonance tongues of phase locked tori, with successive enlargements if one of the tongues in (b) and (c). (©SIAM; Reproduced from Calleja et al. (2017) with permission)
@@
Figure 5.6: (a) An enlargement of Figure 5.5 showing one resonance tongue. (b) A one parameter continuation with $\kappa_{2}=3$ (corresponding to black line in (a) showing values of $u(t-a_{1})$ at each intersection of the orbit with the Poincaré section. (c & d) The phase-locked torus when $(\kappa_{1},\kappa_{2})=(5.79,3)$ (black dot in (a)) with the embedded unstable and stable periodic orbits (c) projected onto $(u(t),u(t-a_{1}),u(t-a_{2}))$-space and (d) its Poincaré trace in the $(u(t-a_{1}),u(t-a_{2}))$ plane. (©SIAM; Reproduced from Calleja et al. (2017) with permission)
@@
Figure 5.7: (a) Detail of the bifurcation diagram from Figure 5.5(a) near to the double-Hopf point $\textit{HH}_{1}$. (b) Inset: Normal form of the bifurcation in the $(\mu_{1},\mu_{2})$ normal form coordinates obtained by computing the normal form coefficients at the Hopf-Hopf point. Main panel: The normal form transformed to the original $(\kappa_{1},\kappa_{2})$ parameters. (©SIAM; Reproduced from Calleja et al. (2017) with permission)

### v2

@@ 10 IVPs with Distributed and Threshold Delays
While it is possible to develop solvers for specific problems (see Langlois et al. (2017); Cassidy et al. (2022)), it is a lot of work to reinvent and recode all the functionality of a sophisticated IVP solver for each new problem. Until a comprehensive numerical RFDE IVP solver is available, it is convenient and time saving to leverage the existing discrete delay software packages to solve problems with distributed delays.
@@ 11 Numerical Continuation and Bifurcation Detection
To apply DDE-Biftool to (2.2),(2.1) we introduce extra dummy delays to discretise the threshold integral. Recall (2.4), which implies that the delay $\tau=\tau(u_{t})$ satisfies
@@
We discretise the time interval $t-\tau_{\textit{max}},t]$ uniformly with a sequence of mesh points
$t=x_{0}>x_{1}>\cdots>x_{N}=t-\tau_{\textit{max}},$
@@
It remains to find $\theta\in[0,1)$ so that $J_{h}(j+\theta)=a$ and hence $\tau_{j}=(j+\theta)a/NV_{\textit{min}}$. To do this we extend the definition of $J_{h}(j)$ to non-integer values of $j$ as illustrated in Figure 4.6. See Gedeon et al. (2022); Wang and Humphries (2022) for further details.
@@ 12 Scalar Threshold Delay Example
Figure 5.1: Bifurcation diagram of (2.2)-(2.1) with $g$ and $V$ both increasing and parameters $\beta=1.4$, $\mu=0.2$, $g^{-}=0.5$, $g^{+}=1$, $\gamma=0.8$, $\theta_{v}=0.5$, $\theta_{g}=a=1$, $v^{-}=0.1$ and $v^{+}=2$. (a) One parameter continuation in $\gamma$ with $m=n=20$. (b) Two-parameter continuations in $m=n$ and $\gamma$ of the fold and the Hopf bifurcations. (Reproduced from Gedeon et al. (2025) under $C$-BY 4.0 license.)
When $g(u)$ and $V(u)$ are both increasing functions, it is possible to have up to five coexisting steady states as illustrated in Figure 5.1(a). In particular, a region of tristability is observed, characterised by three stable steady states separated by two unstable ones. This contrasts sharply with the dynamics of the same operon model with constant state delay, for which at most three steady states are observed when $g$ is an increasing Hill function. If we flip the axes in the Figure 5.1(a) we see that each value of $u>0$ is a steady state for a unique value of $\gamma$, and that the values of $\theta_{g}$ and $\theta_{v}$ organise the locations of the fold bifurcations.
@@
Figure 5.2: One parameter bifurcation diagrams of (2.2)-(2.1) with $g$ constant, and $V$ an increasing function defined by (5.1) with (a) $m=50$ (b) $m=2$ (c) $m=2.25$ and (d) $m=2.35$. Remaining parameters: $\beta=1.4$, $\mu=0.2$, $g^{-}=g^{+}=1$, $\theta_{v}=1$, $a=1$, $v^{-}=0.1$ and $v^{+}=2$. In (b) periodic orbits are represented by the maximum and minimum values of $u$ along the periodic solution profile, and in the other panels by the $L^{2}$-norm (5.2). (Reproduced from Gedeon et al. (2025) under $C$-BY 4.0 license.)
@@
In contrast, in Figure 5.2(b) we plot both $\max u(t)$ and $\min u(t)$ to represent the periodic orbit, which clearly shows the amplitude of the solution. Both representations can be useful on bifurcation diagrams, since all three expressions are equal to the steady state value at a Hopf bifurcation, and thus periodic orbits are seen to emanate from a steady states at a Hopf bifurcation. However, the $L^{2}$-norm (5.2) has the extra property that $\|u\|_{2}\to u_{s}$ as the solution converges to a homoclinic orbit to a saddle steady state $u_{s}$, as seen in both Figures 5.1 and 5.2.
@@ 13 Two Linearly State-Dependent Delays
Figure 5.3: Bifurcation diagrams of (1.22) as $\kappa_{1}$ is varied showing the amplitude and stability of periodic orbits with (a) $\kappa_{2}=1$ (b) $\kappa_{2}=2.3$. In (b) the grey curves denote stable invariant tori. (© AIMS; Reproduced from Humphries et al. (2012) with permission.)
@@
Figure 5.4: Poincaré section $u_{t}(0)=0$ of the quasi periodic torus from Figure 1.5 generated from a single solution of (1.22) with $\kappa_{1}=4.44$ and $\kappa_{2}=3$. (a) Solution segments $u_{t}$ on the Poincaré section. (b) Projection of the solution onto $(u(t),u(t-a_{1}),u(t-a_{2}))$-space (light blue), together with the two-dimensional projection of the Poincaré section. (©SIAM; Reproduced from Calleja et al. (2017) with permission.)
Solutions oscillate about the steady state $u^{*}=0$, so a natural Poincaré section to study the dynamics is $\bigl\{\varphi\in C:\varphi(0)=0$ and $\varphi^{\prime\!}(0)<0\bigr\}$. For finite dimensional dynamical systems the Poincaré section reduces the dimension by one, but $C$ is infinite dimensional, resulting in an infinite dimensional Poincaré section. Plotting function segments on the Poincaré section with $u_{t}(0)=0$ is not very revealing, as seen in Figure 5.4(a).
@@
Figure 5.5: Two-parameter continuation of Hopf (blue), torus (red) and fold (brown) bifurcations in the $(\kappa_{1},\kappa_{2})$-parameter plane show many resonance tongues of phase locked tori, with successive enlargements if one of the tongues in (b) and (c). (©SIAM; Reproduced from Calleja et al. (2017) with permission.)
@@
Figure 5.6: (a) An enlargement of Figure 5.5 showing one resonance tongue. (b) A one parameter continuation with $\kappa_{2}=3$ (corresponding to black line in (a) showing values of $u(t-a_{1})$ at each intersection of the orbit with the Poincaré section. (c & d) The phase-locked torus when $(\kappa_{1},\kappa_{2})=(5.79,3)$ (black dot in (a)) with the embedded unstable and stable periodic orbits (c) projected onto $(u(t),u(t-a_{1}),u(t-a_{2}))$-space and (d) its Poincaré trace in the $(u(t-a_{1}),u(t-a_{2}))$ plane. (©SIAM; Reproduced from Calleja et al. (2017) with permission.)
@@
Figure 5.7: (a) Detail of the bifurcation diagram from Figure 5.5(a) near to the double-Hopf point $\textit{HH}_{1}$. (b) Inset: Normal form of the bifurcation in the $(\mu_{1},\mu_{2})$ normal form coordinates obtained by computing the normal form coefficients at the Hopf-Hopf point. Main panel: The normal form transformed to the original $(\kappa_{1},\kappa_{2})$ parameters. (©SIAM; Reproduced from Calleja et al. (2017) with permission.)